Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 616     Entry time: 28 May 2025, 09:42
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 28.05.2025 
General Announcement:
-	 

CTRL:
- Investigate issue reported last Monday with MID is scheduled for next Monday.

DA:
- Not present.

EEE:
- One topic: motor in CRL in XTD2 found by FXE when doing larger steps caused issues (ZZ seems to be needed); however, with smaller steps it works fine and can be used. 
A ticket is here: https://redmine.xfel.eu/issues/221334#change-751944

VAC:
- Nothing to report.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Adding data space for the last two weeks of operation: a server integration shows an issue and ITDM works on it for the next days.

XRO:
- Not present.

LAS:
- Not present.

UNSYS:
- Nothing to report.

SPB:
- This week still fine; 
- Issues for the Data measured in March with timing drifts and compensation for PP-laser did not work for three weeks. Contact with LAS to be made later today.

FXE:
- Not present.

MID:
- Nothing to report.

HED:
- Not present.

SCS:
- Not present.

SQS:
- Not present.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00